OpenStudy (anonymous):

Find the coordinates for the center for this circle. (x + 3)² + (y - 5)² = 25

OpenStudy (zzr0ck3r):

Given the equation of a circle in the form \((x-a)^2+(y-b)^2=r^2\) the center is \((a,b)\).

OpenStudy (loser66):

I think you know how to do it from the previous one. Why do you post the "almost" same question?

OpenStudy (anonymous):

So (3,5)?

OpenStudy (anonymous):

I'm taking a quiz for my online school and these are the couple I had trouble on.

OpenStudy (anonymous):

Oh excuse me (3, -5)

OpenStudy (zzr0ck3r):

we are not here to take quizes for you

OpenStudy (zzr0ck3r):

quizzes*

OpenStudy (zzr0ck3r):

hint: \(x+3=x-(-3)\)

OpenStudy (anonymous):

Okay would rude. I was just checking to make sure my answer was correct.

OpenStudy (zzr0ck3r):

it is not

OpenStudy (anonymous):

Wow*

OpenStudy (zzr0ck3r):

I am not being rude by telling you the rules:)

OpenStudy (zzr0ck3r):

anyway back to the math...

OpenStudy (zzr0ck3r):

do you see your mistake now?

OpenStudy (anonymous):

I just want to make sure I got the answer right. And yes I do. Thank you.

OpenStudy (plasmataco):

chilllllll, it kinda is rude doe.

OpenStudy (plasmataco):

XD

OpenStudy (zzr0ck3r):

What is your new answer?

OpenStudy (anonymous):

(-3, 5)

OpenStudy (plasmataco):

yup.

OpenStudy (zzr0ck3r):

correct

OpenStudy (plasmataco):

nice job

OpenStudy (anonymous):

Thank you.

OpenStudy (zzr0ck3r):

lol good job @Jcameron